The following data are the measures of the diameters of 36 rivet heads in 1/100 of an inch.

6.72 6.77 6.82 6.70 6.78 6.70 6.62 6.75

6.66 6.66 6.64 6.76 6.73 6.80 6.72 6.76

6.76 6.68 6.66 6.62 6.72 6.76 6.70 6.78

6.76 6.67 6.70 6.72 6.74 6.81 6.79 6.78

6.66 6.76 6.76 6.72

Compute the Arithmetic Mean, variance, standard deviation, Coecient of Variation, Coe-

cient of Skewness and Coecient of Kurtosis..



1
Expert's answer
2022-04-19T03:47:16-0400

First let's sort the data, write appearing frequencies of each value and corresponding probabilities of appearing (the number of different values will be 17):

6.62 - 2, p = 0.117647

6.64 - 1, p = 0.0588235

6.66 - 4, p = 0.235294

6.67 - 1, p = 0.0588235

6.68 - 1, p = 0.0588235

6.7 - 4, p = 0.235294

6.72 - 5, p = 0.294118

6.73 - 1, p = 0.0588235

6.74 - 1, p = 0.0588235

6.75 - 1, p = 0.0588235

6.76 - 7, p = 0.411765

6.77 - 1, p = 0.0588235

6.78 - 3, p = 0.176471

6.79 - 1, p = 0.0588235

6.8 - 1, p = 0.0588235

6.81 - 1, p = 0.0588235

6.82 - 1, p = 0.0588235


Let n = 36


Mean:

"\mu=\frac{1}{n}\sum_{k=1}^nx_k=\\=\frac{1}{36}(6.62\cdot2+6.64+6.66\cdot4+6.67+\\
+6.68+6.7\cdot4+6.72\cdot5+6.73+6.74+\\
+6.75+6.76\cdot7+6.77+6.78\cdot3+6.79+\\
6.8+6.81+6.82)\approx6.73"


Variance:

"\sigma^2=\frac{1}{n-1}\sum_{k=1}^n(x_k-\mu)^2=\\
=\frac{1}{35}(0.11^2\cdot2+0.09^2+0.07^2\cdot4+0.06^2+\\
+0.05^2+0.03^2\cdot4+0.01^2\cdot5+0.01^2+0.02^2+\\
+0.03^2\cdot7+0.04^2+0.05^2\cdot3+0.06^2+\\
+0.07^2+0.08^2+0.09^2)\approx0.002885714\approx0.00289"


Standard deviation:

"\sigma=\sqrt{\sigma^2}\approx0.0537"


Coefficient of variation:

"CV=\frac{\sigma}{\mu}\approx0.008"


Coefficient of skewness:

"b=\frac{1}{n}\sum_{k=1}^n(x_k-\mu)^3/\sigma^3=\\
\frac{1}{36\cdot0.0537^3}((-0.11)^3\cdot2+(-0.09)^3+\\
+(-0.07)^3\cdot4+(-0.06)^3+\\
+(-0.05)^3+(-0.03)^3\cdot4+(-0.01)^3\cdot5+\\
+0.01^3+0.02^3+\\
+0.03^3\cdot7+0.04^3+0.05^3\cdot3+0.06^3+\\
+0.07^3+0.08^3+0.09^3)\approx−18"


Coefficient of Kurtosis:

"K=\frac{n(n+1)}{(n-1)(n-2)(n-3)}\frac{1}{\sigma^4}\sum_{k=1}^n(x_k-\mu)^4=\\
=\frac{36\cdot37}{35\cdot34\cdot33}\frac{1}{0.0537^4}(0.11^4\cdot2+0.09^4+\\
+0.07^4\cdot4+0.06^4+\\
+0.05^4+0.03^4\cdot4+0.01^4\cdot5+0.01^4+0.02^4+\\
+0.03^4\cdot7+0.04^4+0.05^4\cdot3+0.06^4+\\
+0.07^2+0.08^2+0.09^2)\approx2.64"